Please help me solve the following: a 12 liter antifreeze solution is made up of water and alcohol. How much pure alcohol must be added to raise the concentration from 20% to 50%?
Answer by lwsshak3(11628)

Please help me solve the following: a 12 liter antifreeze solution is made up of water and alcohol. How much pure alcohol must be added to raise the concentration from 20% to 50%?
**
let x=amt of pure alcohol to add
20%*12=2.4 liters (amt of pure alcohol in original solution)
2.4/12=20%
(2.4+x)/(12+x)=50%
2.4+x=6+.5x
.5x=3.6
x=7.2
amt of pure alcohol to add=7.2 liters